1. If tn=2n-3, for what value of 'n'is tn=5?
Is it -1,1,4 OR 7??????? I think the answer is 4 or 7????????
2. Which sequence represents an exponential function????
a) {1,6,11,16,21,...}
b) {16,8,4,2,1,...}
c) {2,8,18,32,50,...}
d) {1,8,27,64,125,...} I think it is (b)???
3. Which function represents a geometric sequence?
a) {3*4,3*4^2, 3*4^3,3*4^4,...}
b) {3+4,3+4+4,3+4+4+4,...}
c) {3*4,3*5,3*6,3*7,...}
d) {3+4,3+4^2,3+4^3,...}
Found 2 solutions by stanbon, gonzo: Answer by stanbon(75887) 
 1. If tn=2n-3, for what value of 'n'is tn=5?
Is it -1,1,4 OR 7??????? I think the answer is 4 or 7????????
Correct
--------------------
2. Which sequence represents an exponential function????
a) {1,6,11,16,21,...}
b) {16,8,4,2,1,...}
c) {2,8,18,32,50,...}
d) {1,8,27,64,125,...} I think it is (b)
d = 1^3,3^3,8^3,5^3
Correct
------
3. Which function represents a geometric sequence?
a) {3*4,3*4^2, 3*4^3,3*4^4,...}
b) {3+4,3+4+4,3+4+4+4,...}
c) {3*4,3*5,3*6,3*7,...}
d) {3+4,3+4^2,3+4^3,...}
a has a common ratio of 3
===============================
Cheers,
Stan H.
Answer by gonzo(654) 
 number 1:
answer would be 4.
you simply replace tn with 5 to get 5 = 2n-3.
add 3 to both sides of the equation by 2 to get:
8 = 2n
divide both sides of the equation to get:
n = 4
---
number 2:
not sure about this one.
i don't think it's selection b because that looks like a geometric sequence where r = .5
i don't think it's selection a because that looks like an arithmetic sequence where the value that is added each time is 5.
it's going to be either selection c or selection d.
selection c doesn't appear to rise fast enough to be an exponential function.
8-2 = 6
18-8 = 10
32-18 = 14
50-32 = 18
looks like we have a difference that goes up 4 each time.
a formula that seems to work is:
a[n] = a[n-1] + 2 + 4*(n-1)
when n = 2, a[n] = 2 + 2 + 4*1) = 4 + 4 = 8
when n = 3, a[n] = 8 + 2 + 4*2 = 10 + 8 = 18
when n = 4, a[n] = 18 + 2 + 4*3 = 20 + 12 = 32
etc.
this is not an exponential formula, however, so c is not the answer.
i think it might be selection d.
a formula i think will work there is:
a[n] = n^3
in selection d, a = 1
if n is 1, then a[1] = 1^3 = 1
if n is 2, then a[2] = 2^3 = 8
if n is 3, then a[3] = 3^3 = 27
if n is 4, then a[4] = 4^3 = 64
if n is 5, then a[5] = 5^3 = 125
etc........
since this is an exponential function, then selection d has to be the one.
---
number 3:
geometric sequence would be a[1] * r = a[2] * r = a[3] ..............
looks like it will be selection a.
a[1] = 3
a[2] = 3*4 = 3*4^1
a[3] = (3*4)*4 = 3*4^2
a[4] = (3*4*4)*4 = 3*4^3
etc............
| |
|